OpenStudy (anonymous):

Gracie Shay wants to buy a new Hummer in 5 years. Gracie estimates the cost of the Hummer will be $28,000. If she invests $12,000 now, at a rate of 6 percent compounded semiannually, she: 1.Will have enough money 2.Will have exactly $16,000 3.Will have $18,000 4.Will have $16,126.80 5.None of these

OpenStudy (anonymous):

6/100 * 12,000 = 720 5 YEARS * 2 TIMES IN A YEAR * 720 = 7,200 + 12,000 = $ 19,200 You can see the answer! :)

OpenStudy (anonymous):

I think it's more like: \[$12,000* 1.06^{10} = $21,490.17\] This is because every time you calculate the interest (which is 6%), it includes the interest already added to the principal. And because it's semiannual over 5 years, you calculate interest 10 times.

OpenStudy (anonymous):

\[FV= A/m * [ (1+(i/m ))^ {m*n} - 1 ]/ (i/m)\] FV = Future Value A = Annual lot i = Annual interest rate m = times in a year n = years FV = 720 * [1+6% ]^ 10 - 1 / 6% = 21,490 $ hope it useful .